OpenStudy (anonymous):

the differece of 40 and a natural number is less than 29.5. find the smallest valid natural number

OpenStudy (anonymous):

let the natural number be N then, 40-N=29.5 N=40-29.5 N=10.5 Hence the smallest valid number is 10.5